11123:

At a school carnival, the diameter of the mat of a trampoline is 12 feet and the diameter of its metal frame is 14 feet. What is the length, in feet, of the metal frame that surrounds the trampoline? Use 3.14 for π and round your answer to the nearest tenth. 37.7 feet 44.0 feet 75.4 feet 87.9 feet

11123:

@AnimeGhoul8863

dude:

We need to find circumference (because that is the area that surrounds the trampoline): Circumference = pi \(\times\) diameter \(\color{red}{or}\) \(C=\pi d\) We know: \(\pi=3.14\) and \(d=14\)

11123:

thats not one of the answers

dude:

We have to substitute it into the equation, lol

dude:

\(C=\pi \times d\)

11123:

ohhhh so 44 even?

dude:

Yes

dude:

\(C=3.14\times 14\) \(C=43.96\) which rounds up to 44